sum of the two digit number is 14,if 29 is subtracted then both digits become equal.find the number...
Answer by TutorDelphia(193)

I solved this more as a puzzle than by using algebra.
The possible digits that add up to 14 are
7 and 7, 6 and 8, 5 and 9
So our possible numbers are
77
68
86
59
95
If 29 is subtracted which of these work to make the digits equal?
77-29=48 that doesn't work
68-29=39 that doesn't work
86-29=57 that doesn't work
59-29-30 that doesn't work
95-29=66 yay!